OpenStudy (anonymous):

Solve the equation. Check the solution. -2/x+4 = 4/x+3

OpenStudy (anonymous):

\[\frac{ -2 }{x }+4=\frac{ 4 }{x }+3\] Am i correct ?

OpenStudy (anonymous):

if so, then \[\frac{ 4 }{ x }+\frac{ 2 }{x }=4-3,\frac{ 4+2 }{x }=1,\frac{ 6 }{ x }=1,x=6\]

OpenStudy (anonymous):

Sorry having some issues with the website and pages not loading but here are my answer choices A. -13/6 B.-11 C. -8/3 D. -11/3

OpenStudy (anonymous):

@surjithayer

OpenStudy (whpalmer4):

@surjithayer correctly solved what you wrote, not what you meant. If you are going to write it like that, it needs to be \[-2/(x+4) = 4/(x+3)\]

OpenStudy (whpalmer4):

\[\frac{-2}{x+4}=\frac{4}{x+3}\]Solve this by cross-multiplication: \[-2(x+3) = 4(x+4)\]Expand both sides and solve for \(x\)

OpenStudy (anonymous):

oh my bad :/\[\frac{ -2 }{ x+4 } = \frac{ 4 }{ x+3 }\]

OpenStudy (anonymous):

how do you expand both sides?

OpenStudy (whpalmer4):

\[-2(x+3) = 4(x+4)\]Just apply the distributive property of multiplication:\[a(b+c) = ab + ac\]

OpenStudy (whpalmer4):

So that becomes \[-2*x -2*3 = 4(x+4)\]you do the rest...

OpenStudy (anonymous):

Alright i think i got it thanks

OpenStudy (whpalmer4):

what's your answer?

OpenStudy (anonymous):

I got A

OpenStudy (whpalmer4):

nope. show me your work, let's find the mistake.

OpenStudy (anonymous):

OK so we start with the multiplication right? like with -2*3 = -6?

OpenStudy (whpalmer4):

Yes, so the left side becomes \[-2x-6 = 4(x+4)\]What about the right side?

OpenStudy (anonymous):

x + 16?

OpenStudy (whpalmer4):

no... \[4(x+4) = 4*x + 4*4 = \]

OpenStudy (anonymous):

4x +16?

OpenStudy (whpalmer4):

Right. Want to try solving for \(x\) again? \[-2x-6 = 4x+16\]\[x=\]

OpenStudy (whpalmer4):

Why don't you add \(2x\) to both sides of the equation and tell me what you get?

OpenStudy (anonymous):

x-6 = 6x+16??

OpenStudy (whpalmer4):

No. \[-2x-6= 4x+16\]Add \(2x\) to each side: \[-2x+2x-6 = 4x+16 + 2x\]\[0x - 6 = 6x +16\]\[-6 = 6x+16\]Can you solve that for \(x\)?

OpenStudy (anonymous):

do you add 6 to both sides? Sorry im retaking algebra 2B and i nearly cant remember how to do a lot of things..

OpenStudy (anonymous):

but i really appreciate you helping me

OpenStudy (whpalmer4):

Yes, add 6 to both sides would be one way to go, or subtract 16 from both sides. I'll do both ways: \[-6=6x+16\]\ \[-6+6 = 6x+16+6\]\[0=6x+22\]Subtract 22 from each side\[-22=6x\]Divide each side by 6\]\[\frac{-22}{6} = x\]Reduce that fraction by dividing top and bottom by 2\[-\frac{11}{3} = x\] Or the way I would do it: \[-6=6x+16\]\[-6-16 = 6x+16-16\]\[-22=6x\]and then the same finish

OpenStudy (anonymous):

Alright thanks! :^) I really appreciate the help!

OpenStudy (whpalmer4):

Remember, you can do just about anything to an equation as long as you do the same thing to both sides.